In the spring of​ 2016, the CFO of HTPL Distributing Company decided to distribute a stock dividend to its shareholders.​ Specifically, the CFO proposed that the company pay 0.045 shares of stock to the holders of each share of common stock such that the holder of​ 1,000 shares of stock would receive an additional 45 shares of common stock.

a. If the firm had total net income for the year of $ 9,000,000 and had 22,000,000 shares of common stock outstanding before the stock​ dividend, what was the​ firm's earnings per​ share?

b. After paying the stock​ dividend, what was the​ firm's earnings per​ share?

c. If you owned​ 1,000 shares of stock before the stock​ dividend, how many dollars of earnings did the firm earn from your​ 1,000-share investment? After the stock dividend was​ paid, how many dollars of earnings did the firm earn on your larger share​ holdings? What effect would you expect from the payment of the stock dividend on your total investment in the​ firm?
Answer #1

a. Earnings per share before the stock dividend = $ 9,000,000 / 22,000,000 = $ 0.40909

b. Number of outstanding shares after paying the stock dividend = ( 22,000,000 / 1,000 ) x 45 + 22,000,000 = 22,990,000

Earnings per share after the stock dividend = $ 9,000,000 / 22,990,000 = $ 0.39147

c. If you owned 1,000 shares in the company before the stock dividend, the firm earned $ 0.40,909 x 1,000 = $ 409.09 from your investment.

After the stock dividend, the firm earned 1,045 x 0.39147 = $ 409.09 from your investment.

The total investment in dollar terms does not change.